Determine how much heat is required to bring 750 ml of water from standard room temperature to the boiling point when Standard room temperature is 21° C and the boiling point of water is 99° C.
1
Expert's answer
2020-09-23T09:02:39-0400

Volume of water, "V = 750 ml = 750*0.001*10^{-3} = 7.5*10^{-4} m^3"


Density of water, "\rho = 1000 kg/m^3"

Mass of water, "m = \rho V = 1000 * 7.5*10^{-4} = 0.75 kg"


Heat required, "H = ms(\Delta T) = 0.75 * 4200*78 = 245700 J = 245.7 KJ"

where s is specific heat of water.